OpenStudy (anonymous):

Determine if the series converges or diverges. I am typing my work using equation editor.

OpenStudy (anonymous):

The series is \[\frac{ -3 }{ 12 }+\frac{ 9 }{ 144 }-\frac{ 27 }{ 1728 }+\frac{ 81 }{ 20736 }\] I have determined the series can be represented by the following \[a _{n}= \frac{ (-1)^n }{ 4^n }\] This is where i get stuck. It seems the series converges to "0" since the limit of a equals 0.